Twice last month, Alice rented a car from the same company. She was charged a daily fee, plus an additional charge per mile driven. The first trip lasted 4 days, she drove 380 miles, and the rental cost her $180.20. On her second trip she drove 625 miles in 5 days, and paid $253.75 for the rental. Find the daily fee and the mileage charge.
Answer by mananth(12270)

x Rate per mile
y daily charge
Total memberships
180 x + 4 y = 180.20 .............1

625 x + 5 y = 253.75 .............2
Eliminate y
multiply (1)by -5
Multiply (2) by 4
-900 x -20 y = -901
2500 x + 20 y = 1015
Add the two equations
1600 x = 114
/ 1600
x = 0.07
0.07 *100 = 7 cents/mile
plug value of x in (1)
180 x + 4 y = 180
13 + 4 y = 180
4 y = 180 -13
4 y = 167
y = 42
7 cents per mile & $42 per day